### v1

# Knowledge in multi-robot systems: an interplay of dynamics, computation and communication

Abstract
We show that the hybrid systems perspective of distributed multi-robot systems is compatible with logical models of knowledge already used in distributed computing, and demonstrate its usefulness by deriving sufficient epistemic conditions for exploration and gathering robot tasks to be solvable. We provide a separation of the physical and computational aspects of a robotic system, allowing us to decouple the problems related to each and directly use methods from control theory and distributed computing, fields that are traditionally distant in the literature. Finally, we demonstrate a novel approach for reasoning about the knowledge in multi-robot systems through a principled method of converting a switched hybrid dynamical system into a temporal-epistemic logic model, passing through an abstract state machine representation. This creates space for methods and results to be exchanged across the fields of control theory, distributed computing and temporal-epistemic logic, while reasoning about multi-robot systems.
keywords:
distributed computing, mobile robotics, temporal-epistemic logic, switched system, robot tasks
1 Introduction
In this paper, we provide a connection between the hybrid systems approach to distributed multi-robot systems [34] and temporal epistemic logic via a variation of the runs and systems framework [18], which we construct by defining explicitly a state machine representation of both the system dynamics and the robot protocols. Bridging these fields enables us to combine well-known methods and results from control theory, distributed computing and temporal epistemic logic in a simple and harmonious fashion. This is a non-trivial connection not only because it connects a discrete and a continuous field of research; it also necessarily deviates from the established distributed computing principle that an optimal protocol always exists with respect to information exchange. We further demonstrate the usefulness of our approach by deriving the first ever sufficient epistemic conditions for solving the exploration, surveillance and approximate gathering robot tasks.
The distributed coordination of robots combines problems found in control theory [36, 26] with the problems commonly found in distributed computing [31]. While robotics is traditionally concerned with physical entities that must behave in accordance to their dynamical constraints, the distributed computing perspective arises once the robots must exchange information under varying difficulties associated to their communication capacities [1, 14]. For instance, understanding the impact of synchrony, coordination and message loss are essential for the efficient design and analysis of state-of-the-art multi-robot systems solving various tasks.
In [1] the authors famously showed that exact gathering is impossible by reducing the robot task to the consensus problem in distributed computing and deriving its well understood impossibility borders from there. In essence, this paper puts this reduction on a sound basis by deriving the distributed system implied by the control theoretic model and equipping it with the powerful tools found in epistemic logic.
Epistemic logic [24] is a powerful conceptual tool for reasoning about the uncertainty of agents in a system, both regarding facts and other agents’ epistemic states, making it especially suitable for the modeling of multi-agent systems, including distributed systems [23, 20, 29, 10, 21, 32]. A crucial feature of epistemic logic in multi-agent systems is that it provides a precise and formal representation of higher order epistemic attitudes, such as nested knowledge or nested belief, often crucial for the success of distributed tasks. For instance, coordinated actions are inevitably tied to nested knowledge, i.e., any agent needs to know to some extent what other agents know in order to successfully execute a coordinated behavior [32]. Another interesting feature of epistemic logic is that, when combined with temporal logic, it allows to express properties that must hold (and be known) throughout the execution, without focusing on the specific communication mechanisms leading to such knowledge. Thus, epistemic logic is a universal vehicle for expressing communication models, including those that rely on lack of explicit communication (also known as communication by silence) [20, 21].
The distributed computing community has approached the characterization of tasks in distributed robot systems [19] in various ways, such as according to the system space (continuous or discrete)[3, 13], communication capacities[37] and synchronicity[27]. Model checking using linear temporal logic is also a common tool in this computational perspective of robotics, where it has been used for gathering [16], perpetual exploration [15, 5] and terminating exploration [5].
The control theoretic view on multi-agent systems has been developed by a number of authors, see e.g. the survey [4]. Similarly, logical views on hybrid systems have been considered, see e.g. [33]. As we mentioned also, logics, in particular epistemic logic [23] and temporal-epistemic logic [28] have been developed and applied to multi-agent systems, through in particular epistemic planning [17]. Our approach is the first one to seamlessly integrate three often complementary views: the ones from control theory, distributed systems and temporal epistemic logic.
1.1 Contributions
From the robot abstractions introduced, we define a representation of multi-robot system executions that is pertinent to both the control theory and distributed computing communities. From this common view, we derive a variant of the runs and systems framework wherein atomic propositions represent open sets in a topology, each corresponding to a region of a dense Euclidean space. We then use this epistemic model to provide the first epistemic analysis of key robot tasks, such as exploration, surveillance and approximate gathering, showing that the tasks are solvable under certain (epistemic) conditions. In particular, we prove that the sufficient conditions to solve approximate gathering mimic the sufficient conditions for the stabilizing agreement problem in distributed systems.
This paper is, to the extent of our knowledge, the first framework that connects all three fields comprising control theory, distributed computing, and temporal epistemic logic, in a general setting that enables a comprehensive and integral approach to multi-robot systems.
1.2 Paper organization
The paper is organized as follows. In Section 2 we introduce the basic concepts used for modeling multiple robots: first as a system of ordinary differential equations (Section 2.1), then as state machines (Section 2.2) and finally the correspondence of the two models is shown in Section 2.3. In Section 3, we derive an abstract distributed system from the robot models previously defined. Section 3.1 describes a scheduler that is compatible with both models, followed by matching definitions of system runs using state machines (Section 3.2) and switched systems (Section 3.3). Section 3.4 crucially connects the robotic abstraction to the epistemic model by turning the abstract robot state machine into a suitable epistemic frame. In Section 4, the epistemic frame is enriched into an epistemic model (Section 4.1) , which is then used for the epistemic analysis of robot tasks like simple exploration (Section 4.2), surveillance (Section 4.3) and approximate point gathering (Section 4.4). Finally, concluding thoughts are provided in Section 5.
2 The Robot Model
In this section, we introduce two multi-robot system models with distinct degrees of abstractions and subsequently show one model to be an abstraction of the other: the first is a description of robots as a hybrid dynamical system, close to the traditional approach within the control theory literature [25], the second is a state machine executing look-compute-move cycles, an abstraction commonly found in the distributed computing community [19]. This compatibility allows us to develop a common notion of scheduler for both the dynamical and the computational system perspectives, in Section 3, which will be used to define matching system runs.
2.1 Robots as Dynamical Systems
In the perspective of dynamical systems, robotic systems can modeled via differential equations that capture the evolution of their states over time. Hybrid systems are often used to distinguish the continuous state space of the physical aspects of a robot, from the discrete one of its computations. We first describe the model for a single robot, capable of movement and of computations.
Definition 2.1 (Robot).
A robot is a hybrid system, composed of an ontic state $\mathbf{x}$ and epistemic state $\mathbf{e}$. It has actuators that evolve its ontic state and sensors that perform observations. It has an intelligence and a memory that generate new epistemic states and control commands.
$\dot{\mathbf{x}}(t)$ $=f(\mathbf{x}(t),\mathbf{u}[t])$ (ontic evolution) (1)
$\mathbf{y}(t)$ $=g(\mathbf{x}(t))$ (observation) (2)
$\mathbf{e}[t+1]$ $=\varphi(\mathbf{e}[t],\mathbf{y}[t])$ (epistemic evolution) (3)
$\mathbf{u}[t]$ $=h(\mathbf{e}[t],[t])$ (control) (4)
The ontic evolution $f$ and the observation $g$ occur in continuous time $t$, while the epistemic evolution $\varphi$ and the control $h$ occur in discrete time $[t]$, defined by the staircase function $\left\lfloor t\right\rfloor$.
The ontic state corresponds to the physical state of the robot, while the epistemic state corresponds to the memory state of the robot. The distinction between the ontic and epistemic states allows for the physical and computational problems to be studied separately. As such, a robot’s behavior is given by a (timed) switched system [30], a particular case of hybrid system, as the control $\mathbf{u}[t]$ is possibly discontinuous over time.
Simplifying assumptions can be introduced so that one aspect, epistemic or ontic, is focused over the other. For instance, the epistemic evolution can be rewritten as an output-feedback controller so to recover the traditional control-theoretic point of view [26].
Example 2.2 (State feedback control).
Let a robot be a dynamical system with observer, $\dot{\mathbf{x}}=f(\mathbf{x},\mathbf{u})$ and $\mathbf{y}=g(\mathbf{x})$. An output feedback controller simulates the ontic evolution of the system $\mathbf{x}$ in order to obtain an estimation $\widehat{\mathbf{x}}$ of the state and regulate the input $\mathbf{u}$ of the original system.
$\dot{\widehat{\mathbf{x}}}$ $=f(\widehat{\mathbf{x}},\mathbf{u})$
$\mathbf{u}$ $=h(\widehat{\mathbf{x}},\mathbf{y})$
This controller corresponds to a robot where its epistemic state is an estimation of its ontic state, $\mathbf{e}(t)=\widehat{\mathbf{x}}(t)$, by following the same evolution, $\varphi=f$. The epistemic evolution uses the previous input $\mathbf{u}$ produced by itself for the simulation. The control accesses the observation $\mathbf{y}$, in order to compare with its own, i.e. $\mathbf{u}=h(\widehat{\mathbf{x}},\mathbf{y})=\mathbf{y}-g(\widehat{\mathbf{x}})$.
Alternatively, a point-mass dynamics could be imposed to the ontic evolution in order to obtain a system analogous to distributed computing. Moreover, by assuming a discrete-time dynamics one arrives at a computational model used in complexity theory, see for instance Sections 2.7 and 2.8 in [6].
Example 2.3 (Computing system).
Let a robot be a single computer, with a trivial (constant) ontic evolution $f(\mathbf{x}(t),\mathbf{u}[t])=0$ and an epistemic evolution representing the ’program’. In this context, the literature usually calls them agents.
$\dot{\mathbf{x}}(t)$ $=f(\mathbf{x}(t),\mathbf{u}[t])=0$
$\mathbf{y}(t)$ $=g(\mathbf{x}(t))$
$\mathbf{e}[t+1]$ $=\varphi(\mathbf{e}[t],\mathbf{y}[t])$
$\mathbf{u}[t]$ $=h(\mathbf{e}[t],[t])$
As $\mathbf{x}(t)$ is some constant, $\mathbf{y}(t)$ is also constant. $\varphi$ now only depends on $\mathbf{e}[t]$ and can be seen as the memory of the single computer. The initial external input can be provided via the constant value of $\mathbf{y}(t)$.
Remark 2.4 (From continuous to discrete).
Due to the physical limitations in memory of a physical computing device, we assume that the observation function $g$ has a finite domain $\mathcal{D}(g)$, which corresponds to the digital representation of the continuous quantities it may be capable of measuring. This captures the role of sensors in robotics, such as LIDARs that discretize the environment in distance measurements or temperature sensors that ultimately convert their values to binary via an ADC measuring the voltage drop over a heat sensitive element. The size of $\mathcal{D}(g)$ will depend on how the robots measure their surroundings and derive computable information from it, but we assume that $g$ spans the whole set $\mathcal{D}(g)$, i.e., it is surjective.
A finite and discrete abstraction will be useful in Section 4, where the propositional modal language does not provide quantifiers, restricting to the representation of finite sets of facts. Note that our goal is to derive formulas depending on known facts about the world, therefore reasoning about the robots’ knowledge about facts representing the ontic world in their local memories.
We now consider the modifications required for modeling a multi-robot system where interactions increase in complexity with the added notions of communication and perception among robots. Note that the distinction between communication and perception is a consequence of the distinction between ontic and epistemic states, as both are required to represent the possible information to be exchanged.
Definition 2.5 (Communication and perception of robots).
Let $\mathbf{r}$ and $\mathbf{r}^{\prime}$ be robots. The perception function $\eta$ provides to robot $\mathbf{r}$ an estimation $\widehat{\mathbf{x}}_{\mathbf{r}\mathbf{r}^{\prime}}$ of robot $\mathbf{r}^{\prime}$’s ontic state, and the communication function $\lambda$ provides robot $\mathbf{r}$ with an estimation $\widehat{\mathbf{e}}_{\mathbf{r}\mathbf{r}^{\prime}}$ of the epistemic state of robot $\mathbf{r}^{\prime}$.
$\widehat{\mathbf{x}}_{\mathbf{r}\mathbf{r}^{\prime}}(t)$ $=\eta_{\mathbf{r}}(\mathbf{x}_{\mathbf{r}^{\prime}}(t))$ (perception) (5)
$\widehat{\mathbf{e}}_{\mathbf{r}\mathbf{r}^{\prime}}[t]$ $=\lambda_{\mathbf{r}}(\mathbf{e}_{\mathbf{r}^{\prime}}[t])$ (communication) (6)
Communication and perception complement the observation step with information about the other robots, which in turn is used by the epistemic evolution.
$\begin{split}\mathbf{y}_{\mathbf{r}}(t)=g(\mathbf{x}(t),\{\eta_{\mathbf{r}}(\mathbf{x}_{\mathbf{r}^{\prime}}(t))\},\{\lambda_{\mathbf{r}}(\mathbf{e}_{\mathbf{r}^{\prime}}[t])\})\hskip 30.00005pt\text{(multi-robot observation)}\end{split}$ (7)
Note that Definition 2.5 places communication and perception as two types of interaction that are processed together by the epistemic evolution of the robot in Equation 3. This understanding will match the behavior of observation operations in the perspective of robots as computing systems.
2.2 Robots as Mobile Computing Systems
By discerning the physical dynamics of the robot from the dynamics of its computations, it becomes evident that the epistemic evolution matches the role of an algorithm, where the epistemic state is the memory being operated on. Now we will explore how the robots described in Definition 2.1 can be abstracted as mobile computing systems executing look-compute-move cycles in a state machine.
Definition 2.6 (look-compute-move cycle [19]).
The look-compute-move model decomposes the behavior of robots in an environment as cycles that execute three active steps and one inactive step. The environment consists of the universe in which the robots move (for example a graph or a continuous region of space), and some adversarial behavior influencing the robots’ actions. The adversarial definition includes at least a scheduler deciding on which robot takes the next step in the look-compute-move cycle.
  1. 1.
    Look. The robot makes use of it sensors to retrieve information from the environment. A snapshot displays the physical state of the other robots, possibly limited by a visibility radius.
  2. 2.
    Compute. The information obtained in the Look step is used to locally compute the succeeding memory state. All robots execute the same deterministic algorithm.
  3. 3.
    Move. The robot changes its physical state (based on its current memory state), which most commonly consists of its position in the environment.
  4. 4.
    Wait. The robot does not perform any action, but remains observable.
Uncertainty in the behavior of each step, such as the accuracy of the information or the actual moved distance, are possible choices for adversarial behavior in the LCM model. Some variations are depicted in Appendix A.
Note that the physical dynamics of a robot are only relevant when interacting with the environment, by updating (move) or by retrieving (look) its state. The computational aspects of a robot are encapsulated by the compute step, that matches the processing of the acquired information and the generation of commands to be executed.
The simplest models of robots running lcm cycles are not capable of interaction besides observing the existence and physical state of each other via look operations. Therefore, we will consider the enhanced model of luminous robots, equipped with programmable lights for communication and memory. Note that this is a parallel requirement to the communication and perception introduced in Definition 2.5.
Definition 2.7 ($\mathcal{LUMI}$ model [12]).
The LUMInous robot model consists of a group of homogeneous robots operating lcm cycles. They are modeled as points in an environment $\mathcal{E}$ that operate without central control. Each robot possesses a set of external lights that can be modified in order to express some information for communication, as they can be observed by other robots, or for memory, as they are preserved throughout rounds. The physical state, including the lights, are updated during a move step. During look operations, each robot obtains a snapshot of the environment $\mathcal{E}$, which provides the physical states, including the lights, of all other robots. In the weakest version, robots with the same lights are indistinguishable and have no identity. They are disoriented and possess no global reference frame.
Remark 2.8.
Note that the communication made available through the $\mathcal{LUMI}$ model is indirect, as any message provided by a robot must be actively observed by another one for it to be received. Here, perception and communication are combined in a single interaction. The outgoing communication is entirely comprised by the actions that the robot can make, i.e., set the lights according to its new state. Symmetrically, the incoming communication is entirely described by the observations made by a robot.
Now, we describe the state machine executed by a robot operating under the $\mathcal{LUMI}$ model. We consider that a robot $\mathbf{r}$ exists in an environment $\mathcal{E}$ and executes $\mathcal{A}$ during the compute step of the lcm cycle. The evolution of $\mathcal{E}$ comprehends both of the move and look steps, where the ontic updates of all robots are executed and the possible observations retrieved.
Definition 2.9 (Robot state machine).
A robot $\mathbf{r}$ executes the state machine
$\mathcal{A}=\langle S_{\mathbf{e}_{\mathbf{r}}},S_{\mathbf{y}_{\mathbf{r}}},S_{\mathbf{u}_{\mathbf{r}}},g_{\mathbf{r}},\varphi_{\mathbf{r}},h_{\mathbf{r}}\rangle$
where
  • $S_{\mathbf{e}_{\mathbf{r}}}$ is a finite set of possible memory states;
  • $S_{\mathbf{y}_{\mathbf{r}}}$ is a finite set of possible observations obtained from the environment $\mathcal{E}$;
  • $S_{\mathbf{u}_{\mathbf{r}}}$ is a set of actions that can be executed in the environment $\mathcal{E}$;
  • $g:\mathcal{E}\to S_{\mathbf{y}_{\mathbf{r}}}$ is the function that retrieves an observation from the current environment $\mathcal{E}$;
  • $\varphi:S_{\mathbf{e}_{\mathbf{r}}}\times S_{\mathbf{y}_{\mathbf{r}}}\to S_{\mathbf{e}_{\mathbf{r}}}$ is a function taking an epistemic state and an observation to a new epistemic state;
  • $h:S_{\mathbf{e}_{\mathbf{r}}}\to S_{\mathbf{u}_{\mathbf{r}}}$ is a function that generates an action given an epistemic state.
The environment itself is dynamic and therefore also a state machine. It provides the observations used by the robots during their execution of $\mathcal{A}$ and evolves according to the robots’ actions and some possible adversarial factor. It abstracts the ontic state of all robots as a single entity, with which all of the robot state machines interact.
Definition 2.10 (Environment state machine).
The environment evolves according to the state machine
$\mathcal{E}=\langle S_{\mathcal{E}},I_{\mathcal{B}},\{S_{\mathbf{u}_{\mathbf{r}}}\}_{\mathbf{r}\in\Pi},\{S_{\mathbf{y}_{\mathbf{r}}}\}_{\mathbf{r}\in\Pi},F_{\mathcal{E}},G_{\mathcal{E}}\rangle$
where
  • $S_{\mathcal{E}}$ is a set of possible environment states;
  • $I_{\mathcal{B}}$ is a set of possible choices made by the adversary;
  • $\{S_{\mathbf{u}_{\mathbf{r}}}\}_{\mathbf{r}\in\Pi}$ is a collection of possible robot actions;
  • $\{S_{\mathbf{y}_{\mathbf{r}}}\}_{\mathbf{r}\in\Pi}$ is a collection of possible observations;
  • $F_{\mathcal{E}}:S_{\mathcal{E}}\times\prod_{\mathbf{r}\in\Pi}S_{\mathbf{u}_{\mathbf{r}}}\times I_{\mathcal{B}}\to S_{\mathcal{E}}$ is the function generating a new environment state from the latest actions and the adversarial interference;
  • $G_{\mathcal{E}}:S_{\mathcal{E}}\times I_{\mathcal{B}}\to\prod_{\mathbf{r}\in\Pi}S_{\mathbf{y}_{\mathbf{r}}}$ is the function that generates observations from the current environment state and possible adversarial interference.
This depiction allows us to focus on the computations executed locally, while the environment abstracts all ontic states and updates into a single entity, with which all robots’ state machines interact. The look and move are split into distinct functions that interface the environment and robot: one of preparation of the data ($h_{\mathbf{r}}$ and $G_{\mathcal{E}}$), and one of usage of the data, $F_{\mathcal{E}}$ and $g_{\mathbf{r}}$. This split will facilitate stating the correspondence between computational and dynamical perspectives in Section 2.3.
This abstraction will provide the algorithmic perspective of a robotic system essential to obtaining a principled representation of the knowledge available in Section 4.
2.3 Correspondence of the Dynamical and Computational Perspectives
We now have the language to state the compatibility of the dynamical robot model described in Section 2.1 and the computational model described in Section 2.2. Let $\mathbf{r}$ be a robot modeled as a hybrid system according to Definition 2.1 with communication and perception as in Definition 2.5. Let $\mathbf{r}^{\prime}$ be a luminous robot, as in Definition 2.7 executing lcm cycles as in Definition 2.9, and let $\mathcal{E}$ be the environment in which $\mathbf{r}^{\prime}$ executes, as in Definition 2.10. We have that:
Proposition 2.11 (Correspondence of discretized state spaces).
The following functions that map the space of $\mathbf{r}$ to that of $\mathbf{r}^{\prime}$ are surjective.
  1. (i)
    $\Phi_{\mathbf{e}}:\mathbf{e}_{\mathbf{r}}\to S_{\mathbf{e}_{\mathbf{r}^{\prime}}}$, from epistemic states to memory states;
  2. (ii)
    $\Phi_{\mathbf{y}}:\mathbf{y}_{\mathbf{r}}\to S_{\mathbf{y}_{\mathbf{r}^{\prime}}}$, from observations to discrete observations, as per Remark 2.4;
  3. (iii)
    $\Phi_{\mathbf{u}}:\mathbf{u}_{\mathbf{r}}\to S_{\mathbf{u}_{\mathbf{r}^{\prime}}}$, from a discrete control to discrete actions;
  4. (iv)
    $\Phi_{\mathcal{E}}:\mathbf{x}_{\mathbf{r}}\to S_{\mathcal{E}_{\mathbf{r}^{\prime}}}$, from ontic states to environment states.
Proposition 2.12 (Correspondence of dynamics).
For any execution of $\mathbf{r}$ over a time interval $[t,t+\Delta]$ corresponding to one lcm cycle of $\mathbf{r}^{\prime}$, we have that:
  1. (i)
    $\varphi_{\mathbf{r}}:\mathbf{e}_{\mathbf{r}}[t]\times\mathbf{y}_{\mathbf{r}}[t]\to\mathbf{e}_{\mathbf{r}}[t+1]$ corresponds to $\varphi_{\mathbf{r}^{\prime}}:S_{\mathbf{e}_{\mathbf{r}^{\prime}}}\times S_{\mathbf{y}_{\mathbf{r}^{\prime}}}\to S_{\mathbf{e}_{\mathbf{r}^{\prime}}}$. Also, $h:\mathbf{e}_{\mathbf{r}}[t]\to\mathbf{u}_{\mathbf{r}}[t]$ corresponds to $h_{\mathbf{r}^{\prime}}:S_{\mathbf{e}_{\mathbf{r}^{\prime}}\to S_{\mathbf{u}_{\mathbf{r}^{\prime}}}}$. That is, the epistemic and control functions match the robot update.
  2. (ii)
    $f_{\mathbf{r}}(\cdot)$ corresponds to $F_{\mathcal{E}}(\cdot)$. That is, the ontic evolution corresponds to the environment update.
  3. (iii)
    $g_{\mathbf{r}}(\cdot)$ corresponds to $g_{\mathbf{r}^{\prime}}(G_{\mathcal{E}}(\cdot))$. That is, the observation functions match.
Proposition 2.13 (Correspondence of interactions).
The communication $\lambda_{\mathbf{r}_{1}\mathbf{r}_{2}}(\cdot)$ and perception $\eta_{\mathbf{r}_{1}\mathbf{r}_{2}}(\cdot)$ functions, in Equation 7 of Definition 2.5, alongside the general environment perception $g_{\mathbf{r}}$ correspond to the components of the function $G_{\mathcal{E}}(\cdot)$ generating observations from the environment in Definition 2.10. That is, the communication and perception functions correspond to the observation generation in the environment state machine.
3 Distributed Multi-Robot Systems and Epistemic Frames
In this section, we will define the construction of system runs, and support it by a general notion of scheduler, applicable to both the dynamical and computational perspectives provided in Section 2. A general epistemic frame for multi-robot tasks will then be introduced using the system runs as its building block. The epistemic frames will be later specialized, in Section 4, to runs and systems models fitting different robot tasks.
3.1 Scheduler: Global and Local Time
We introduce the notion of time path in order to connect the local executions of robots in a distributed system to a global time variable. We consider that robots execute their protocols in a linearly ordered time set, either +:=$\mathbb{R}_{+}:=\{x\in\mathbb{R}\;|\;x\geqslant 0\}$, or $\mathbb{N}$. Individually, each robot executes sequentially [^1] This is for simplicity and it should be possible to relax the assumption for a concurrent cycle at the cost of extra bookkeeping of how each step uses the information from the others to execute. the cycle of look, compute and move steps. We consider that cycles start with the move step, corresponding to an initial empty look and compute steps, so to match with the dynamical system definition.
The robots’ order of activation is defined by a scheduler, which chooses the robot that will execute the next step of its cycle. Common schedulers found in the literature are the synchronous, where all steps happens simultaneously, semi-synchronous, where the synchronization is guaranteed at every full cycle, and asynchronous, where there is no shared notion of time. Check [19] for further details on the taxonomy of schedulers. Any scheduler that is not fully synchronous may lead to global time that differs from the local clocks a robot may be equipped with. This distinction is made precise with the notion of a time path, matching the role of a global time.
[…]
Definition 3.1 (Time path).
[…]
Furthermore, we suppose that $p$ is rectified in the sense that the length of the path [9] from $p(0)$ to $p(t)$ for all $t\in[0,T]$ has length $t$, using the infinity distance defined by $d_{\infty}((x_{1},\ldots,x_{n}),(y_{1},\ldots,y_{n}))=max(|x_{1}-y_{1}|,|x_{2}-y_{2}|,\ldots,|x_{n}-y_{n})$.
Given a time path $p$, the local time of robot $\mathbf{r}$ is given by $t_{\mathbf{r}}$ or $[t]_{\mathbf{r}}$, depending on if it is continuous or discrete.
Such a time path defines the relative progress of every robot at play, where a scheduler corresponds to a family of time paths. For instance, a synchronous scheduler can only correspond to a diagonal straight line, where all operations are aligned, while any path representing a semi-synchronous scheduler needs to cross all compute steps simultaneously.
Remark 3.2 (Notation for time).
It is important to note the different notations used for time. We use $t$ for the continuous time and $[t]$ for discrete time, as per Definition 2.1. A subscript with the robot identifier is used to depict local time, as in $t_{\mathbf{r}}$ or $[t]_{\mathbf{r}}$, and its absence represents global time.
We should note that, depending on the modeling choices, robots may or may not have access to a local clock that allows them to estimate or approximate the global time. Two possible executions of a multi-robot system composed of two robots is illustrated in Figure 1. In the case of instantaneous moves as in [1], because the behavior of robots would be undefined at points where a look operation (or event) would happen at the same time as a move operation, one should forbid points (represented as small squares in Figure 1) at the intersection of move and look events.
$[t]_{2}$$[t]_{1}$$\textsc{m}_{0}$$\textsc{l}_{0}$$\textsc{c}_{0}$$\textsc{m}_{1}$$\textsc{l}_{1}$$\textsc{c}_{1}$$\textsc{m}_{0}$$\textsc{l}_{0}$$\textsc{c}_{0}$$\textsc{m}_{1}$$\textsc{l}_{1}$$\textsc{c}_{1}$$\textsc{m}_{2}$$\textsc{l}_{2}$$\textsc{c}_{2}$$\textsc{m}_{3}$$\textsc{l}_{3}$$\textsc{c}_{3}$
Figure 1: Two global time paths, blue and red, represent different possible executions of robots with local times $[t]_{1}$ and $[t]_{2}$. Each robot executes the sequence of actions (M) update ontic state, (L) observe environment and other robots, (C) update epistemic state. Each moment the time path crosses a dotted line, the associated robot executes the corresponding action. The actions are indexed by the cycle they belong to.
Note that communication scheduling is comprised in the current definition of the scheduler. Once the communication between robots is defined as part of the perception in the $\mathcal{LUMI}$ model or in Definition 2.5, a “delayed” message can be simplified to a robot not having executed the operation that would retrieve that information yet.
3.2 Executions of Distributed Multi-Robot Systems
Remarkably, the definitions of global and local times associated to a time path correspond precisely to the clock of the environment and the robot, respectively, in Definitions 2.9 and 2.10. This allows us to describe the robot executions, looking only at their local computations.
Definition 3.3 (Execution).
Let $\mathbf{r}\in\Pi$ be a robot, with an epistemic evolution specified by Definition 2.9. Let a pair $\langle\mathbf{e}_{\mathbf{r}},\mathbf{y}_{\mathbf{r}}\rangle$ be a robot configuration. We say that an execution of $\mathbf{r}$ is a sequence $\{\langle\mathbf{e}_{\mathbf{r}}[t]_{\mathbf{r}},\mathbf{y}_{\mathbf{r}}[t]_{\mathbf{r}}\rangle\}$ indexed by its local time $[t]_{\mathbf{r}}\in N$, where $\mathbf{e}_{\mathbf{r}}[t]_{\mathbf{r}}\in S_{\mathbf{r}}$, $\mathbf{y}_{\mathbf{r}}[t]_{\mathbf{r}}\in S_{\mathbf{y}}$, and $\varphi_{\mathbf{r}}(\mathbf{e}_{\mathbf{r}}[t]_{\mathbf{r}},\mathbf{y}_{\mathbf{r}}[t]_{\mathbf{r}})=\mathbf{e}_{\mathbf{r}}[t+1]_{\mathbf{r}}$.
We can extend the notion of a robot execution to a system run, arriving at a global description of the system. We need first to define global robot configurations, which will take the role of possible worlds in the epistemic frame later described in Section 3.4.
Definition 3.4 (Global robot configuration).
A global robot configuration is tuple $\mathcal{C}=\langle\{\mathbf{e}_{\mathbf{r}}\}_{\mathbf{r}\in\Pi},\{\mathbf{y}_{\mathbf{r}}\}_{\mathbf{r}\in\Pi},[t]\rangle$ consisting of all local robot configurations at a global computation time $[t]$. $\mathcal{C}_{\mathbf{r}}$ is the restriction of a global configuration $\mathcal{C}$ to the robot $\mathbf{r}$.
Definition 3.5 (System run).
A system run $\sigma$ from time 00 to $T$ is a sequence of global robot configurations $\{\mathcal{C}[t]\}_{[t]\in[0,T]}$, where, for any $t\in[0,T]$:
  • there exists a subset $\varnothing\neq P(t)\subseteq\Pi$ of participating robots.
  • we have
    $\mathcal{C}_{\mathbf{r}}[t+1]=\\ \begin{cases}\varphi_{\mathbf{r}}(\mathcal{C}[t]_{\mathbf{r}})&\textrm{ if }\mathbf{r}\in P(t)\\ \mathcal{C}[t]_{\mathbf{r}}&\textrm{ if }\mathbf{r}\notin P(t)\end{cases}$
We denote the set of all possible runs $\mathcal{I}$, representing all the possible evolutions of the system solely based on the robot’s behavior. Note that this definition of run abstracts away the physical updates, in similar fashion to how the robot state machine represents the robots’ computing capacities. We effectively focus on the epistemic states, as the environment encompasses the interactions and constraints obtained from the ontic evolution and observations across all robots. The executions correspond to the individual epistemic evolution of a single robot, while the runs represent the epistemic evolution of all the robots at compatible cuts.
3.3 Solutions of Distributed Multi-Robot System
We attest now the uniform treatment of time provided by the scheduler by bridging the understanding of system runs of robots as dynamical systems (Section 2.1) similarly to that of computing systems (Section 2.2), followed with a temporal-epistemic model pertinent to both perspectives in Section 4.
A distributed multi-robot system becomes an amalgamation of the individual robots obeying similar dynamics and epistemic update rules. The description of the global system is simplified by combining the state variables and observations as follows.
For simplicity, we assume that all robots live in a hypercube bounded by the $[0,1]$ interval $X\subseteq[0,1]^{k}$, with $k\geq 1$, along with its standard Euclidean metric, i.e. $d(\bar{x},\bar{y})=\sqrt{\sum_{i=1}^{k}(x_{i}-y_{i})^{2}}$, where $\bar{x}=(x_{1},\ldots,x_{k})$ and $\bar{y}=(y_{1},\ldots,y_{k})$. We consider only the case of homogeneous multi-robot systems, equipped with identical capacities and behavior, but note that the described model does not limit to such case.
Definition 3.6 (Distributed multi-robot system).
A distributed multi-robot system consists of $N$ robots, capable of observing the environment and interacting with the other robots through perception and communication. For a given time path $p$, its evolution is given by:
$\dot{\mathbf{X}}(t)$ $=F_{p}(\mathbf{X}(t),\mathbf{U}[t])$ (global ontic evolution) (8)
$\mathbf{Y}(t)$ $=G(\mathbf{X}(t),\eta(\mathbf{X}(t)),\lambda(\mathbf{E}[t]))$ (multi-robot observation) (9)
$\mathbf{E}[t+1]$ $=\Phi(\mathbf{E}[t],\mathbf{Y}(t))$ (global epistemic evolution) (10)
$\mathbf{U}[t]$ $=H(\mathbf{E}[t],t)$ (global control) (11)
This definition is a consequence of the evolution of single robots, as seen in Definitions 2.1 and 2.5. Changing variables in Equation 1, so that to consider the states of each robots at global time $t$ for some time path $p$, $\mathbf{X}(t)=\mathbf{x}(p_{i}(t))$, we get:
$\dot{\mathbf{X}}(t)$ $=\langle\dot{p},f(\mathbf{X}(t),\mathbf{u}[t])\rangle$
[…]
Equations (2), (3) and (4) now translate into, in terms of state $\mathbf{X}(t)$
$\mathbf{Y}(t)$ $=G(\mathbf{X}(t))$ (observation) (12)
$\mathbf{E}[t+1]$ $=\Phi(\mathbf{E}[t],\mathbf{Y}(t))$ (epistemic evolution) (13)
$\mathbf{U}[t]$ $=H(\mathbf{E}[t],t)$ (control) (14)
where $G=g$, $\Phi=\varphi$ and $H(\mathbf{E}[t],t)=h(\mathbf{E}[t],[t]_{\mathbf{r}})$.
The (arbitrary) switched system of Definition 3.6 has a notion of solution, which will be used later as the dynamical system version of “runs”. For this, we need to see our switched system as a more general differential inclusion:
Definition 3.7 (Solutions to differential inclusions[2]).
Consider the general differential inclusion
$\dot{x}\in\mathbb{F}(x)$ (15)
[…]
In general, there can be many solutions to a differential inclusion. Throughout the section we note $S_{\mathbb{F}}(x_{0})$ the set of all (absolutely continuous) solutions to the Equation 15.
[…]
Definition 3.8 (Marchaud map[2]).
The set-valued map $\mathbb{F}:\mathbb{R}^{n}\leadsto\mathbb{R}^{n}$ is a Marchaud map if $\mathbb{F}$ is upper semicontinuous with compact convex values and linear growth (that is, there is a constant $c>0$ such that $|\mathbb{F}(x)|:=sup\{|y|\ |\ y\in\mathbb{F}(x)\}\leqslant c(1+|x|)$, for every $x$).
We know from [2] that when $\mathbb{F}$ is a Marchaud map, then the inclusion (15) has a solution such that $x(t_{0})=x_{0}$ (for all $x_{0}$) and for a sufficiently small time interval $[t_{0},t_{0}+\varepsilon)$, $\varepsilon>0$. Global existence, for all $t\in\mbox{$\mathbb{R}$}$ can be shown provided $\mathbb{F}$ does not allow “blow-up” ($\|x(t)\|\rightarrow\infty$ as $t\rightarrow t^{*}$ for a finite $t^{*}$).
In that case, we can produce system runs from any solution $x$ of switched system (treated as a differential inclusion) of Equation (8) defined on time $[0,T]$, as in Definition 3.5 as follows. The system run is $\{\mathcal{C}(t)\}_{t\in[0,T]}$ is given by
$\mathcal{C}(t+1)_{\mathbf{r}}=\\ \begin{cases}G_{\mathbf{r}}(x(t),\eta(x(t)),\lambda(\mathbf{E}[t]))&\textrm{ if }\mathbf{r}\in P(t)\\ \mathcal{C}_{\mathbf{r}}(t)&\textrm{ if }\mathbf{r}\notin P(t)\end{cases}$
where $\mathbf{E}[t]$ is defined as in Equation (13).
3.4 System Runs and Epistemic Frames
The remaining of this paper will focus on the computational aspect of the distributed multi-robot systems described in Section 3.3, i.e., on their epistemic states and on the transitions between epistemic states. The main tool that we are going to use to describe the epistemic evolution of robotic systems are epistemic frames, that we build from the system runs of Definition 3.5:
Definition 3.9 (Epistemic Frame).
Given a set of robots $\Pi$ and the set of all possible runs $\Sigma$, an epistemic frame =$\mathcal{F}=\langle\{\mathcal{C}\}_{\mathcal{C}\in\Sigma},\{\sim_{\mathbf{r}}\}_{\mathbf{r}\in\Pi}\rangle$ is composed of a set of global robot configurations $\{\mathcal{C}\}_{\mathcal{C}\in\Sigma}$ (Definition 3.4) and a collection of indistinguishability relations $\{\sim_{\mathbf{r}}\}_{\mathbf{r}\in\Pi}$, where:
  • Each global robot configuration, or point, is constituted by a pair $(\sigma,[t])$ of system run $\sigma$ (Definition 3.5) and the corresponding global timestamp $[t]$ (Definition 3.1), which is relative to the run $\sigma\in\Sigma$;
  • the collection of indistinguishability relations $\sim_{\mathbf{r}}$ indexed by robot $\mathbf{r}\in\Pi$ are computed based on the equivalence of the local state: $(\sigma,[t])\sim_{\mathbf{r}}(\sigma^{\prime},[t^{\prime}])$ iff $e_{\mathbf{r}}^{\mathcal{C}}=e_{\mathbf{r}}^{\mathcal{C}^{\prime}}$ where $\sigma[t]=\mathcal{C}$, i.e., if the epistemic state $e_{\mathbf{r}}$ of $\mathbf{r}$ in two different global robot configurations is the same.
Remark 3.10.
The structure of the epistemic frame, and in particular its indistinguishability relations, depends heavily on the assumptions of the robot model (Table 1). To give some examples, if one assumes non-rigid movement then the robots will have greater uncertainty concerning their current location; visibility influences how much information can be gained during a look step; asynchronicity allows indistinguisbility to range among points with exact same global state but different global time, and so on.
4 Robot Tasks and Temporal Epistemic Logic
In this section we use the obtained epistemic frames to reason epistemically about robot tasks. We start by formalizing exploration tasks, and then we extend and refine this framework so as to accommodate the specifics of surveillance and approximate gathering tasks.
In robot tasks, we assume that robots are able to sense a portion of the environment in which they are embedded, typically represented as a topological space. We call the “observable” portion of the environment the exploration space, which we assume to be a compact topological space, denoted by $(X,\tau)$. Recall that a pair $(X,\tau)$ is a topological space if $X$ is a set, $\tau\subseteq 2^{X}$, $\varnothing\in\tau$, $X\in\tau$, and $\tau$ is closed under arbitrary unions, and finite intersections. $X$ is usually called the “space”, and $\tau$ is the collection of open sets. Recall that $(X,\tau)$ is compact iff for any $S\subseteq 2^{\tau}$ such that $\bigcup_{U\in S}U=X$, then there exists a finite subset $S^{\prime}\subseteq S$ such that $\bigcup_{U\in S^{\prime}}U=X$. Recall that a set $V\subseteq X$ is closed if $X\setminus V\in\tau$. For simplicity and illustrative purposes, we will assume that the exploration space is a compact normalized Euclidean space [^2] The Heine-Borel theorem states that in Euclidean spaces, compactness is equivalent to the space being bounded and closed in $\mathbb{R}^{n}$., i.e. $X\subseteq[0,1]^{n}$ with the standard Euclidean metric. Recall that in a metric space, a set $U\subseteq X$ is open iff for any $x\in U$, there exists an $\varepsilon>0$ such that $B_{\varepsilon}(x):=\{y\in X\;|\;d(x,y)<\varepsilon\}\subseteq U$.
Since we are interested in determining whether or not a region has been explored, we represent statements about exploration, such as “region $U$ has been explored”, using propositions: for every open set $U$ of $\tau$, we associate a unique proposition $\text{exp}(U)$, representing the fact that region $U$ is explored. $\text{exp}(X)$ expresses the fact that the whole space $X$ is explored, and $\text{exp}(\varnothing)$ represents that no portion of $X$ is explored. Because of the natural correlation between proposition and explored space, we refer to $\text{P}_{X}$ as the set of exploration statements.
Exploration statements satisfy some important properties: $V\supseteq U$ implies that $\text{exp}(V)\rightarrow\text{exp}(U)$. Thus, $(\text{P}_{X},\rightarrow)$ is a poset, induced by the containment of open sets. Furthermore, since $(X,\tau)$ is a topological space, and thus closed under union, then $(\text{P}_{X},\rightarrow)$ is a join-lattice. This property allows us to express exploration in a compact way, namely, by considering the largest explored region instead of considering separately all the possible smaller regions that have been explored. For instance, a robot a need only communicate to another robot b, the largest possible region that it has explored.
4.1 A Temporal Epistemic Logic for Exploration
One main advantage of epistemic frames is that we can interpret indistinguishability relations as capturing the robots’ knowledge. Intuitively, a robot $\mathbf{r}$ knows a formula $\psi$ at $(\sigma,[t])$ iff such a formula holds at all points that are indistinguishable from $(\sigma,[t])$ to $\mathbf{r}$. In particular, because we defined indistinguishability between points on equivalence of robots’ local states, an S5 modal logic [^3] In an S5 modal logic, the knowledge operator satisfies factivity ($K_{a}\psi\rightarrow\psi$), positive introspection ($K_{a}\psi\rightarrow K_{a}K_{a}\psi$) and negative introspection ($\neg K_{a}\psi\rightarrow K_{a}\neg K_{a}\psi$). is the natural candidate. This notion of knowledge can be extended to groups, such as distributed knowledge via the indistinguishability relation $\sim_{\mathcal{D}_{A}}:=\bigcap_{\mathbf{r}\in A}\sim_{\mathbf{r}}$ consisting of the intersection of indistinguishabilities. It is easy to verify that $\sim_{\mathcal{D}_{A}}$ is also an equivalence relation. Intuitively, distributed knowledge represents the “collective knowledge” that is spread across all robots in $A$. Another important notion of group knowledge is mutual knowledge, which corresponds to everybody in a group knows a given formula $\psi$.
An epistemic frame can be enriched to an epistemic model by adding an evaluation function that assigns a truth value to atoms at each point of the model [^4] Since we focus mainly on exploration tasks, we restrict our attention to exploration statements, $\text{P}_{X}$. Nevertheless, note that the set of propositions may be chosen otherwise, and which one suits the best depends on the task at hand..
Formally, an epistemic model $\mathcal{M}$ is a tuple of an epistemic frame $\mathcal{F}$ and an evaluation function $\pi:\text{P}_{X}\rightarrow 2^{\{\mathcal{C}\}_{\mathcal{C}\in\Sigma}}$, where $\mathcal{C}$ is a global robot configuration, and $\text{P}_{X}$ is the collection of exploration statements. In our particular case, the function $\pi$ simply maps a proposition $\text{exp}(U)$ corresponding to a region $U$, to all the global configurations where $U$ has been jointly explored (with or without communication) by the robots. In line with the runs and systems framework, we call our epistemic model interpreted system, denoted by $\mathcal{I}$.
Remark 4.1.
Note that, in the interpreted system usually found in the literature, a ”full-information“ protocol is used as a simple protocol that is comparable to the best obtainable one. Such protocol is not possible to be found for the case of distributed multi-robot systems, as robots necessarily need to make mutually exclusive choices, such as which direction to go. Therefore we fix a protocol, in the function $\varphi$ of Definition 3.6, and consider that indistinguishability in our runs ranges over the possible adversarial behavior of the environment.
The following grammar defines temporal epistemic logic formulas for exploration:
Definition 4.2 (Language $\mathcal{L}_{exp}$).
The exploration language is defined by the following grammar:
$\varphi::=\text{exp}(U)\>|\>\neg\varphi\>|\>(\varphi\wedge\varphi)\>|\>K_{\mathbf{r}}\varphi\>|\>D_{A}\varphi\>|\>\Diamond\varphi$
where $\text{exp}(U)\in\text{P}_{X}$, $\mathbf{r}\in\Pi$, and $A\subseteq\Pi$.
In this grammar, $\text{exp}(U)$ is an exploration statement, $K_{\mathbf{r}}\varphi$ represents “$\mathbf{r}$ knows $\varphi$”, $D_{A}\varphi$ stands for “ group A has distributed knowledge of $\varphi$”, and $\Diamond\varphi$ reads “eventually $\varphi$”.
Definition 4.3 (Semantics of $\mathcal{L}_{exp}$).
Let $\mathcal{I}$ be a set of runs, with its respective set of configuration-time pairs $\mathcal{W}$, an evaluation function $\pi:\text{P}_{X}\rightarrow 2^{\mathcal{W}}$, a run $\sigma\in\mathcal{I}$, and a robot $\mathbf{r}\in\Pi$, :
  • $\mathcal{I},(\sigma,[t])\vDash\text{exp}(U)\text{ iff }\text{exp}(U)\in\pi(\text{exp}(U))$
  • $\mathcal{I},(\sigma,[t])\vDash\neg\varphi\text{ iff }\mathcal{I},(\sigma,[t])\not\vDash\varphi$
  • $\mathcal{I},(\sigma,[t])\vDash(\varphi\wedge\psi)\text{ iff }\mathcal{I},(\sigma,[t])\vDash\varphi\text{ and }\mathcal{I},(\sigma,[t])\vDash\psi$
  • $\mathcal{I},(\sigma,[t])\vDash K_{\mathbf{r}}\varphi\text{ iff }\mathcal{I},(\sigma^{\prime},[t]^{\prime})\vDash\varphi\text{ for all }(\sigma^{\prime},[t]^{\prime})\in\mathcal{W}\text{ s.t. }(\sigma,[t])\underset{\mathbf{r}}{\sim}(\sigma^{\prime},[t]^{\prime})$
  • $\mathcal{I},(\sigma,[t])\vDash D_{A}\varphi\text{ iff }\mathcal{I},(\sigma^{\prime},[t]^{\prime})\vDash\varphi\text{ for all }(\sigma^{\prime},[t]^{\prime})\in\mathcal{W}\text{ s.t. }(\sigma,[t])\underset{{\mathcal{D}_{A}}}{\sim}(\sigma^{\prime},[t]^{\prime})$
  • $\mathcal{I},(\sigma,[t])\vDash\Diamond\varphi$ iff $\mathcal{I},(\sigma,[t]^{\prime})\vDash\varphi$ for some $[t]^{\prime}\geqslant[t]$.
We further set $\Box\varphi:=\neg\Diamond\neg\varphi$ and $\varphi\vee\psi:=\neg(\neg\varphi\wedge\neg\psi)$. We also define mutual knowledge $E_{\Pi}\varphi$ as the conjunction of the knowledge of robots in $\Pi$: $E_{\Pi}\varphi:=\bigwedge_{\mathbf{r}\in\Pi}K_{\mathbf{r}}\varphi$.
Parenthesis are omitted in favor of readability whenever precedence is not ambiguous.
A formula $\varphi$ is valid in $\mathcal{I}$ if, for all $(\sigma,[t])\in\mathcal{W}$, $\mathcal{I},(\sigma,[t])\vDash\varphi$. If $\varphi$ is valid in $\mathcal{I}$, we use the notation $\mathcal{I}\vDash\varphi$.
Remark 4.4.
If an robot knows that a certain region $\text{exp}(U)$ has been explored, formally $K_{\mathbf{r}}\text{exp}(U)$, it implies that all subsets of that region are also known by that robot, i.e., $K_{\mathbf{r}}\text{exp}(V)$ with $V\subseteq U$ and $V$ open. This follows directly from the compactness of the exploration space.
Having established the temporal epistemic framework, it is evident that the space will be explored at a run $\sigma\in\mathcal{I}$ iff $\mathcal{I},\sigma\vDash\Diamond\text{exp}(X)$.
Although this condition seems simple enough to achieve, a more interesting part of the analysis lies within the different termination conditions for the exploration problem. In what follows, we introduce the epistemic characterization of the simple exploration robot task and different termination conditions for it.
Definition 4.5 (Simple Robot Exploration).
We say that a system is consistent with a simple exploration system if it satisfies:
  • Exploration agency:
    $\mathcal{I}\vDash\text{exp}(U)\rightarrow D_{\Pi}\text{exp}(U)$ (16)
  • Exploration independence:
    $\mathcal{I}\vDash D_{A}\text{exp}(U)\rightarrow\bigwedge_{\mathbf{r}\in A}K_{\mathbf{r}}exp(V_{\mathbf{r}})\textrm{ such that }\bigcup_{\mathbf{r}\in A}V_{\mathbf{r}}=U$ (17)
  • Stable exploration statements:
    For any region $U$
    $\mathcal{I}\vDash\bigwedge_{\text{exp}(U)\in\mathcal{L}_{exp}^{-}}\text{exp}(U)\rightarrow\Box\text{exp}(U)$ (18)
  • Stable exploration knowledge (perfect recall):
    For any $\mathbf{r}\in\Pi$, and any region $U$
    $\mathcal{I}\vDash K_{\mathbf{r}}\text{exp}(U)\rightarrow\Box K_{\mathbf{r}}\text{exp}(U)$ (19)
Exploration agency claims that an atom is true, meaning that a region is explored, because it has been explored by at least one robot in the group.
Exploration independence means that whatever exploration atom the robots know distributively, is given by the union of the exploration atoms known by those robots. In other words, exploration (distributed) knowledge can be acquired either by direct observation, or robot communication and not by other inferences about the system.
Stable exploration statements states that once a region has been explored, it will be remain forever explored. Stable exploration knowledge says that once an exploration statement is known, it will be known forever by that robot.
The following definitions formally state different notions of termination that may be considered for the exploration task, and possibly for robot tasks in general.
Definition 4.6 (Termination Conditions).
Given an interpreted system $\mathcal{I}$ that is consistent with simple exploration, a set of robots $\Pi$, with an exploration space $X$, and its respective exploration statements $\text{P}_{X}$, we define the following termination conditions:
  • Parallel Termination: We say that $\mathcal{I}$ is consistent with parallel termination iff $\mathcal{I}\vDash\Diamond D_{\Pi}\text{exp}(X)$.
  • Selfish Termination: We say that $\mathcal{I}$ is consistent with selfish termination iff $\mathcal{I}\vDash\Diamond E_{\Pi}\text{exp}(X)$.
  • Cooperative Termination: We say that $\mathcal{I}$ is consistent with cooperative termination iff $\mathcal{I}\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{exp}(X)$.
Intuitively, the parallel termination condition corresponds to a model where a global control or observer is able to terminate the task whenever it is completed, the selfish termination condition corresponds to a setting where it is sufficient for each robot to be aware that the exploration is complete in order to terminate, and cooperative termination corresponds to a setting where a robot may only terminate whenever it is certain that eventually all other processes will be aware that the space exploration is complete.
In the following, we will characterize parallel termination in terms of a very weak condition, which we call exploration liveness. Intuitively, exploration liveness means that any point $x\in X$ of the exploration space $X$ will eventually be explored. Note that, since we assume that robot exploration space consists only of open sets of $X$, the exploration liveness condition reduces to the existence of a (possibly infinite) open cover of $X$ that consists of explored regions, i.e., there must exist a collection $\mathcal{U}$ of open sets, such that $\bigcup_{U\in\mathcal{U}}U=X$, and eventually each $U$ is explored.
Definition 4.7 (Exploration Liveness).
We say that a run $\sigma$ satisfies exploration liveness iff there exists a collection of open sets, $\mathcal{U}$, such that $\bigcup_{U\in\mathcal{U}}U=X$, and for any $U\in\mathcal{U}$, $\mathcal{I},\sigma\vDash\Diamond\text{exp}(U)$. Respectively, we say that a model satisfies exploration liveness iff all of its runs satisfy exploration liveness.
Theorem 4.8 (Exploration Liveness Necessity).
Exploration liveness is necessary for simple exploration.
Proof 4.9.
Note by contraposition that if there is a run $\sigma$, such that no open cover $\mathcal{U}$ of eventually explored regions exists. Naturally, we may consider $\mathcal{U}^{\prime}=\{U\;|\;\mathcal{I},\sigma\vDash\Diamond\text{exp}(U)\}$. Since $\mathcal{U}^{\prime}$ is not an open cover of $X$, then $\bigcup_{U\in\mathcal{U}^{\prime}}U\neq X$. Therefore, there is a point $x\in X,x\notin\bigcup_{U\in\mathcal{U}^{\prime}}U$. Thus, by definition of $\mathcal{U}^{\prime}$, $x$ is never explored, and consequently $X$ is not explored in $\sigma$.
Theorem 4.10 (Exploration Liveness Sufficiency).
Let $\mathcal{I}$ be a model with a compact exploration space $X$. If $\mathcal{I}$ satisfies exploration liveness, then $\mathcal{I}\vDash\Diamond\text{exp}(X)$.
Proof 4.11.
Consider an arbitrary run $\sigma$ of $\mathcal{I}$, from Definition 4.7, there is an open cover $\mathcal{U}$ such that for any $U\in\mathcal{U}$, $\mathcal{I},\sigma\vDash\Diamond\text{exp}(U)$. Since $X$ is compact and $\mathcal{U}$ is an open cover of $X$ then there exists a finite subcover $\mathcal{U}^{\prime}=\{V_{1},\ldots,V_{k}\}\subseteq\mathcal{U}$, such that $\bigcup_{i=1}^{k}V_{i}=X$. Also, since $\mathcal{I},\sigma\vDash\Diamond\text{exp}(V_{i})$, then for any $i\in[1,k]$, then for each $V_{i}$ there is a time $[t_{i}]$ such that $\mathcal{I},\sigma,[t_{i}]\vDash\text{exp}(V_{i})$. Simply consider $[t_{\textrm{max}}]=\mbox{}_{\textrm{max}}\{[t_{i}]\}$. Since exploration statements are stable facts, then $\mathcal{I},\sigma,[t_{\textrm{max}}]\vDash\bigwedge_{i=1}^{k}\text{exp}(V_{i})$. Therefore $\mathcal{I},\sigma,[t_{\textrm{max}}]\vDash\text{exp}(X)$. Thus $\mathcal{I},\sigma\vDash\Diamond\text{exp}(X)$. Since $\sigma$ is an arbitrary run of $\mathcal{I}$, then it follows that $\mathcal{I}\vDash\Diamond\text{exp}(X)$.
Note that Theorem 4.8, Theorem 4.10, and exploration agency from Definition 4.5 imply that whenever $X$ is compact, then exploration liveness is necessary and sufficient for parallel termination. However this is not necessarily true in non-compact spaces. For instance, consider a single robot exploring the open interval $(0,1)\subset\mathbb{R}$. Consider a scenario where at each step $[t]$ the robot has explored the region $(0,1-(1/2)^{[t]})$. It is easy to verify that in this case, the exploration liveness condition is satisfied, nevertheless $\Diamond\text{exp}(X)$ never holds, since at any given time, there is still a portion of the space that remains to be explored, namely $[1-(1/2)^{[t]},1)$.
4.2 The Exploration Task in the LCM Model
We first consider a $\mathcal{LUMI}$ robot system, as defined in Definition 2.7. Recall that robots communicate by displaying information via a set of lights that is assumed to remain active at all times, and that in the full-visibility case is observable at any time by an active robot in the look phase of its LCM cycle.
Definition 4.12 (Exploration Flooding).
We say that the communication of a robot system is exploration flooding iff at each phase of the LCM cycle, each robot tries to communicate the largest region it knows to be explored to the rest of the robots.
Lemma 4.13.
Consider a model $\mathcal{I}$ with an exploration space $X$, from a $\mathcal{LUMI}$ robot system with full visibility and exploration flooding, and $\text{exp}(U)$ an exploration statement, an arbitrary run $\sigma$ from $\mathcal{I}$, and an arbitrary robot $\mathbf{r}\in\Pi$. If $\mathcal{I},\sigma\vDash\Diamond K_{\mathbf{r}}\text{exp}(U)$, then $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\text{exp}(U)$.
Proof 4.14.
Since $\mathcal{I},\sigma\vDash\Diamond K_{\mathbf{r}}\text{exp}(U)$, then there exists a time $[t_{0}]$ such that $\mathcal{I},\sigma,[t_{0}]\vDash K_{\mathbf{r}}\text{exp}(U)$. Since robots have perfect recall, and exploration statements are stable, $\mathcal{I},\sigma,[t]\vDash K_{\mathbf{r}}\text{exp}(U)$ for any $[t]\geqslant[t_{0}]$.
Let $\mathbf{r}^{\prime}\in\Pi$ be an arbitrary robot. Since we don’t consider crash failures, there is a time $[t]_{\mathbf{r}^{\prime}}\geqslant[t_{0}]$ when $\mathbf{r}^{\prime}$ is able to read the lights of robot $\mathbf{r}$. Let $U^{\prime}$ be the maximal region that $\mathbf{r}$ knows to be explored at time $[t]_{\mathbf{r}^{\prime}}$. Since the robot system is exploration flooding, then robot $\mathbf{r}$ is communicating $\text{exp}(U^{\prime})$ through its lights. Therefore, robot $\mathbf{r}^{\prime}$ learns $\text{exp}(U^{\prime})$, i.e. $K_{\mathbf{r}^{\prime}}\text{exp}(U^{\prime})$. However, recall that $U\subseteq U^{\prime}$. Thus $\mathcal{I},\sigma,[t]_{\mathbf{r}^{\prime}}\vDash K_{\mathbf{r}^{\prime}}\text{exp}(U)$.
Since $\Pi$ is finite, then there is a $[t_{\textrm{max}}]$ such that $\mathcal{I},\sigma,[t_{\textrm{max}}]K_{\mathbf{r}^{\prime\prime}}\text{exp}(U)$ for any robot $\mathbf{r}^{\prime\prime}\in\Pi$. Finally, $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\text{exp}(U)$.
Theorem 4.15.
Consider an interpreted system $\mathcal{I}$ with a compact exploration space $X$ such that $\mathcal{I}$ satisfies the exploration liveness condition, and such that the robot system is exploration flooding. Then $\mathcal{I}$ satisfies the cooperative termination property from Definition 4.6, namely, $\mathcal{I}\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{exp}(X)$
Proof 4.16.
Since $X$ is compact and satisfies exploration liveness, then by Theorem 4.8, then $\mathcal{I}\vDash\Diamond\text{exp}(X)$. By exploration agency, then $\mathcal{I}\vDash\Diamond D_{\Pi}\text{exp}(X)$. Let $\sigma$ be a fixed run of $\mathcal{I}$, then there exists a time $[t_{0}]$ such that $\mathcal{I},(\sigma,[t_{0}])\vDash D_{\Pi}\text{exp}(X)$. Consider for each robot $\mathbf{r}\in\Pi$, a region $U_{\mathbf{r}}:=\bigcup\{U\subseteq X\;|\;\mathcal{I},\sigma,[t_{0}]\vDash K_{\mathbf{r}}\text{exp}(U)\}$. Note that $K_{\mathbf{r}}\text{exp}(U_{\mathbf{r}})$, and $U_{\mathbf{r}}$ is the maximal region that robot $\mathbf{r}$ knows to be explored. From exploration independenceDefinition 4.5), it follows that $\bigcup_{i=1}^{k}U_{i}=X$.
From Lemma 4.13, it holds that $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}U_{\mathbf{r}}$. Therefore, $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\bigcup_{i=1}^{n}U_{i}$. This shows that $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\text{exp}(X)$.
Let $[t_{1}]$ be such that $\mathcal{I},\sigma,[t_{1}]\vDash E_{\Pi}\text{exp}(X)$. In particular consider an arbitrary robot $\mathbf{r}^{\prime}\in\Pi$. Since $\mathcal{I},\sigma,[t_{1}]\vDash K_{\mathbf{r}}\text{exp}(X)$, then by definition of knowledge, in any other run $\sigma^{\prime}$ and any other time $[t^{\prime}]$ that is indistinguishable from $\sigma$ at time $[t_{1}]$ to $\mathbf{r}^{\prime}$, $\text{exp}(X)$ holds. From Lemma 4.13, it follows that $\mathcal{I},\sigma^{\prime},[t^{\prime}]\vDash\Diamond E_{\Pi}\text{exp}(X)$. Thus, from the definition of the knowledge modality, $\mathcal{I},\sigma,[t_{1}]\vDash K_{\mathbf{r}^{\prime}}\Diamond E_{\Pi}\text{exp}(X)$. Since $\mathbf{r}^{\prime}$ is arbitrary, then $\mathcal{I},\sigma,[t_{1}]\vDash E_{\Pi}\Diamond E_{\Pi}\text{exp}(X)$. Therefore $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{exp}(X)$. Therefore, $\mathcal{I}\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{exp}(X)$.
Corollary 4.17.
All termination conditions are equivalent for a $\mathcal{LUMI}$ robot system with exploration flooding, and exploration fairness.
Proof 4.18.
Note that Cooperative Termination $\Rightarrow$ Selfish termination $\Rightarrow$ Parallel Termination. Finally note that Theorem 4.15 shows that Parallel Termination $\Rightarrow$ Cooperative Termination, which completes the equivalence cycle.
We can extend our result to robot systems with arbitrary communication, nevertheless, we still need to define a liveness condition for the communication regarding exploration.
Definition 4.19 (Communication Liveness).
We say that a model $\mathcal{I}$ with an exploration space $X$ satisfies the communication liveness condition if for any run $\sigma$ of $\mathcal{I}$, any time $[t]$, any robot $\mathbf{r}$, and any exploration statement $p_{u}$, $\mathcal{I},(\sigma,[t])\vDash K_{\mathbf{r}}\text{exp}(U)\Rightarrow\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\text{exp}(U)$.
Theorem 4.20.
Consider a model $\mathcal{I}$ with a compact exploration space $X$ such that $\mathcal{I}$ satisfies the exploration liveness and communication liveness conditions, and such that the robot system is exploration flooding. Then $\mathcal{I}$ satisfies the cooperative termination property from Definition 4.6, namely, $\mathcal{I}\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{exp}(X)$
Proof 4.21.
Note that we can reuse the proof of Theorem 4.15, but using Definition 4.19 instead of Lemma 4.13.
4.3 The Surveillance Task
In the surveillance task, we consider that an external agent might invade the robots space at a given time frame. If such an intruder happens to invade the space, then the robots should be able to detect it. In the following, we will show that the surveillance task can be formulated as a variant of the exploration task.
We will consider a normalized time frame for detection, represented by the closed interval $[0,1]\subset\mathbb{R}$. As with exploration, we consider a compact space $X$, which we will call our surveillance space. Note that the product space of our surveillance space with the normalized time frame, $X\times[0,1]$ is a compact topological space with the product topology, which we will call the surveillance cylinder.
Note that each exploration region observed by the robots in the surveillance also induces an exploration region in the surveillance cylinder by simply taking the time of the observations into account. For instance, if a robot explores a region $U$ of the surveillance space at a time $t$, then it already has explored $U\times t$ within the surveillance cylinder. Furthermore, if the intruder’s speed is known to be bounded by a factor of $f$, then the exploration of $U$ at time $t_{1}$, implies that that at the next time $t_{2}$ at which the robot is active, $(U\setminus B_{f\cdot t}(\partial U)),t)_{t\in[t_{1},t_{2}]}$ is also explored, where $B_{f\cdot t}(\partial U)$ is the ball of radius $f\cdot t$ around the boundary of $U$. This comes from the fact that if the intruder was not observed inside $U$ at time $t_{1}$, then the farthest that the intruder can reach at a time $t\in[t_{1},t_{2}]$ is given by the ball of radius $f\cdot t$ around the boundary of $U$.
For the surveillance task, we will consider a set of propositional atoms consisting of exploration statements of the surveillance cylinder $X\times[0,1]$, namely $\text{P}_{X\times[0,1]}$. Each of these exploration statements represent that a part of the surveillance cylinder was explored and no intruder was found. In addition to $\text{P}_{X\times[0,1]}$, we also consider two additional propositional atoms, FOUND, and SECURE. Thus, we can define a set of propositional atoms for surveillance, denoted by $\text{P}_{Surv(X)}:=\text{P}_{X\times[0,1]}\cup\{\text{FOUND},\text{SECURE}\}$. Furthermore, we consider an epistemic temporal language for surveillance in the same way as we did for exploration, but with $\text{P}_{Surv(X)}$ instead of $\text{P}_{X}$.
Definition 4.22 (Language $\mathcal{L}_{surv}$).
The surveillance language is defined by the following grammar:
$\varphi::=p\>|\>\neg\varphi\>|\>(\varphi\wedge\varphi)\>|\>K_{\mathbf{r}}\varphi\>|\>D_{A}\varphi\>|\>\Diamond\varphi$
where $p\in\text{P}_{Surv(X)}$, $\mathbf{r}\in\Pi$, and $A\subseteq\Pi$.
The semantics of $\mathcal{L}_{surv}$ is analogous to the one of $\mathcal{L}_{exp}$ in Definition 4.3. Note that the surveillance space is considered to be secure iff there does not exist a path $\xi:[0,1]\rightarrow X$ such that the trace of $\xi$. $tr(\xi):=\{(\xi(t),t)\;|\;t\in[0,1]\}$ is not explored. A naive approach to secure the surveillance cylinder would be to consider the exploration on $X\times[0,1]$. However this is largely impractical, and often unfeasible, as it is equivalent to observing the space $X$ at all times during the time interval $[0,1]$. Instead, it is sufficient to explore a "cut" or section of the surveillance cylinder. For instance, if the surveillance space $X$ was explored completely at a fixed time time, i.e., if the set $X\times\{t\}$ from the surveillance cylinder is explored, then the path of any intruder would intersect $X\times\{t\}$. More generally, if there is a manifold in the surveillance cylinder that has been observed, and can be expressed as the trace of a continuous function $f:X\rightarrow[0,1]$; $\text{tr}(f):=\{(x,f(x))\;|\;x\in X\}$.
Lemma 4.23 (Curve catching lemma).
Let $\alpha:[0,1]\rightarrow X$ be an arbitrary curve in $X$, and $f:X\rightarrow[0,1]$ a continuous function. Then, the trace of $\alpha$, $\text{tr}(\alpha):=\{(\alpha(t),t)\;|\;t\in[0,1]\}$ and the trace of $f$, $\text{tr}(f):=\{(x,f(x))\;|\;x\in X\}$ intersect.
Proof 4.24.
Simply note that $f\circ\alpha:[0,1]\rightarrow[0,1]$ is a continuous function on $[0,1]$. Thus by the Brouwer’s fixed-point theorem, there must exist a $t_{0}$ such that $f\circ\alpha(t_{0})=t_{0}$. Let $\alpha(t_{0})=y$. Thus $(y,t_{0})\in\text{tr}(\alpha)$ and by definition $(\alpha(t_{0}),f(\alpha(t_{0})))\in\text{tr}(f)$. Therefore $(y,t_{0})\in\text{tr}(\alpha)\cap\text{tr}(f)$.
Intuitively, this lemma shows that continuous manifolds on $X$ work as a "catching net" that is able to capture the path of any invader. Therefore, in addition to the conditions required for simple exploration ( Definition 4.5), we also consider the following conditions:
Definition 4.25 (Found and Secure Conditions).
Detection Agency Consider an arbitrary run $\sigma$ in a robot system, and an intruder path $\alpha:[0,1]\rightarrow X$. We say that FOUND holds at a time $t$ iff either there exists a robot $p_{i}$ whose observation region at time $t$ intersects $\alpha$, or there is a time $t^{\prime}<t$ such that FOUND holds at time $t^{\prime}$.
Observation Correctness Consider an arbitrary run $\sigma$, and an intruder path $\alpha:[0,1]\rightarrow X$. Let $U$ be a region of the surveillance cylinder $X\times[0,1]$ such that $U\cap\text{tr}(\alpha)\neq\varnothing$. Then $\neg\text{exp}(U)$.
Securement Let $f:X\rightarrow[0,1]$ be any continuous function. Let $U$ be a region of the surveillance cylinder $X\times[0,1]$, if $\text{tr}(f)\subseteq U$, then $\text{exp}(U)\rightarrow\text{SECURE}$.
Completion condition $\mathcal{I}\vDash(\text{SECURE}\implies\Box\neg\text{FOUND})\wedge(\text{FOUND}\implies\Box\neg\text{SECURE})$.
Detection agency means that the FOUND atom only holds if an intruder has already been detected by a robot. Observation correctness means that a region in the surveillance cylinder is not secure if an intruder is moving through a space-time point in the region, i.e., exploration regions in this context refer only to regions where no intruder has been found. Securement captures the idea from Lemma 4.23 that shows that it is sufficient to explore the trace of a continuous manifold in $X$. Completion condition states that if part of the surveillance cylinder was explored then the intruder was not found (and never will), and if the intruder was found then the surveillance cylinder was not secure (and never will be). Also note that if the trace of the manifold has even a single discontinuity point, then this may be enough for an intruder to evade detection.
We provide the following definition for the surveillance robot task (without termination):
Definition 4.26 (Surveillance).
Consider a robot system and a model $\mathcal{I}$ consistent with Definition 4.5 with respect to its surveillance cylinder, that satisfies Definition 4.25. Let $\sigma$ be a run of $\mathcal{I}$. We say that the model solves the surveillance task if $\mathcal{I},\sigma\vDash\Diamond\text{FOUND}\vee\Diamond\text{SECURE}$.
Note that by Completion condition (Definition 4.25) SECURE and FOUND atoms are mutually exclusive in a run.
As with the simple exploration task, we can consider multiple termination conditions for the surveillance task.
Definition 4.27 (Termination for surveillance).
Parallel Termination $\mathcal{I},\sigma\vDash\Diamond D_{\Pi}\text{FOUND}\vee\Diamond D_{\Pi}\text{SECURE}$
Selfish Termination $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\text{FOUND}\vee\Diamond E_{\Pi}\text{SECURE}$
Cooperative Termination $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{FOUND}\vee\Diamond E_{\Pi}\Diamond E_{\Pi}\text{SECURE}$
Since Lemma 4.23 shows that exploring a continuous manifold of the surveillance cylinder is enough to detect an intruder, it is also enough to fully secure the space, thus we can replace the exploration liveness condition to a manifold instance.
Definition 4.28 (Manifold liveness).
We say that run $\sigma$ of a robot system has the manifold liveness condition iff there exists a continuous function $f_{\sigma}X\rightarrow[0,1]$ such that $\sigma$ satisfies the exploration liveness condition for $\text{tr}(f_{\sigma})$.
Theorem 4.29 (Manifold liveness sufficiency).
Consider a robot system that satisfies the simple exploration conditions, Definition 4.5, for the surveillance cylinder $X\times[0,1]$, the surveillance conditions, Definition 4.25, and the manifold surveillance liveness condition, Definition 4.28. Then the surveillance task is solved at $\sigma$.
Proof 4.30.
Follows from Lemma 4.23, and Theorem 4.10.
Theorem 4.31.
Consider a model $\mathcal{I}$ with a compact surveillance space $X$ that satisfies the communication liveness, the manifold liveness condition, the simple exploration conditions for the surveillance cone $X\times[0,1]$, the surveillance conditions, and the exploration flooding with respect to the surveillance cylinder. Then $\mathcal{I}$ satisfies the cooperative termination property.
Proof 4.32.
Let $\sigma$ be a run of $\mathcal{I}$, from Theorem 4.29, $\mathcal{I},\sigma\vDash\Diamond\text{FOUND}\vee\Diamond\text{SECURE}$. Recall that by Completion condition (Definition 4.25) SECURE and FOUND are mutually exclusive in a run. First, let us assume that $\mathcal{I},\sigma\vDash\Diamond\text{FOUND}$. From Definition 4.25, there must exist a robot $\mathbf{r}_{i}$ such that finds an intruder at a time $t_{i}$. Since $\mathbf{r}_{i}$ detects the intruder, it holds that $\mathcal{I},\sigma,t_{i}\vDash K_{\mathbf{r}_{i}}\text{FOUND}$. From Definition 4.19, it holds that $\mathcal{I}\sigma\vDash\Diamond E_{\Pi}\text{FOUND}$. In particular, in all runs $\sigma^{\prime}$ that are indistinguishable to $\mathbf{r}_{i}$, $\mathcal{I},\sigma^{\prime}\vDash\Diamond E_{\Pi}\text{FOUND}$. Therefore, $\mathcal{I},\sigma,t_{i}\vDash K_{\mathbf{r}_{i}}\Diamond E_{\Pi}\text{FOUND}$. Since $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\text{FOUND}$, then we can repeat the previous argument for any other robot. Therefore $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{FOUND}$.
Now, assume that $\mathcal{I},\sigma\vDash\Diamond\text{SECURE}$ holds. Then, it follows that no intruder was found. From Definition 4.28, it holds that there exists a continuous manifold $f_{\sigma}:X\rightarrow[0,1]$, such that $\mathcal{I},\sigma\vDash\Diamond p_{\text{tr}(f_{\sigma})}$. Note that the conditions for Theorem 4.20 hold, therefore, $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}p_{\text{tr}(f_{\sigma})}$. From the securement condition from Definition 4.25, it follows that $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{SECURE}$.
Corollary 4.33.
In particular for the $\mathcal{LUMI}$ model with full visibility, if the robot system satisfies the manifold liveness, then it solves the surveillance task with cooperative termination
4.4 Approximate Point Gathering Task
In the approximate gathering task, we require robots to eventually meet up in the same space.
Instead of gathering at a precise point, robots need to ensure that their position and all other robots positions’ are contained in a specific open set, called the rendezvous area. To this extent, we need to formulate statements depending on the robots current position, thus extending the previously introduced epistemic language with positional atoms, representing the position where the robot currently sits. Furthermore, we now explicitly require the set of possible observed positions $\text{P}_{g(X)}$ (i.e., a subset of the domain of $g$ in Definition 2.1) to be finite over the exploration space $X$, as set in Remark 2.4. $\text{P}_{g(X)}$ is a finite subset of our space $\text{P}_{g(X)}\subseteq X$, such that $|\text{P}_{g(X)}|\in\mbox{$\mathbb{N}$}$, where $\vec{x}\in\text{P}_{X}^{\Pi}$ denotes a vector with a position for each robot. If $\text{P}_{g(X)}$ does not cover the whole space, then there are places where a robot simply cannot go.
We denote its positional atoms by $pos_{\mathbf{r}}(x)$ meaning that robot $\mathbf{r}$ is at some actual position $\mathbf{x}$ that is discretized to $x$. Note that, as $g$ is surjective, we might have $g_{\mathbf{r}}(\mathbf{x})=g_{\mathbf{r}}(\mathbf{x}^{\prime})$ even though $\mathbf{x}\neq\mathbf{x}^{\prime}$, meaning two different actual robot positions result in the same position for that robot. By construction, we assume that robots know their own position: $pos_{\mathbf{r}}(x)\leftrightarrow K_{\mathbf{r}}pos_{\mathbf{r}}(x)$. We accordingly extend the semantics by $\pi(pos_{\mathbf{r}}(x))$ to contain any point $(\sigma,[t])$ where robot $\mathbf{r}$ believes to be at the position $x$. We further equip the language with facts about the inclusion of points in open sets in the exploration space $\in_{x,U}$ denoting that and abbreviated as $x\in U$ collected in the set $\text{P}_{\tau}$. Note that, since the position and subset atoms are parameterized by points and open sets, we can form expressions connecting them. For example $pos_{\mathbf{r}}(x)\wedge(x\in U)$ holding, denotes that robot $\mathbf{r}$ believes to be position $x$ which is completely contained in area $U$.
Additionally we identify by $init(pos_{\mathbf{r}}(x))$ the initial canonical position of robot $\mathbf{r}$, meaning its starting position, and extend its semantics via $\pi(init(pos_{\mathbf{r}}(x))=\{(\sigma,t)\;|\;(\sigma,0)\vDash pos_{\mathbf{r}}(x)\}$. Lastly, we define a finite set of exploration atoms to denote possible rendezvous areas $\text{P}_{rz(X)}\in 2^{X}$ such that $\forall U,V\in\text{P}_{rz(X)}:U\cap V=\varnothing$ and $|\text{P}_{rz(X)}|\in\mbox{$\mathbb{N}$}$. The approximate point gathering is solved if all robots’ positions are contained in the same rendezvous area forever, assuming each robot starts in one rendezvous area. By the previous reasoning, the robots also know their initial starting position. Naturally, the robots need not do anything when all their initial canonical position are in the same rendezvous area. Otherwise they need to agree on one rendezvous area and move to there. They are allowed to change the rendezvous area finitely often, but eventually need to stick to one forever. We require it to be impossible for any robot to intersect two rendezvous areas at the same time, i.e., $pos_{r}(x)\wedge(x\in U)\implies(\bigwedge_{U\neq V\in\text{P}_{rz(X)}}\neg(x\in V))$. This resembles the notion of taking a unique choice, where we forbid a robot to choose two areas simultaneously.
The following grammar defines temporal epistemic logic formulas for approximate point gathering.
Definition 4.34 (Language $\mathcal{L}_{gather}$).
The gathering language is defined by the following grammar:
$\varphi::=p\>|\>\neg\varphi\>|\>(\varphi\wedge\varphi)\>|\>K_{r}\varphi\>|\>D_{A}\varphi\>|\>\Diamond\varphi$
where $p\in\text{P}_{\tau}\cup\text{P}_{g(X)}\cup\text{P}_{rz(X)}$, $\mathbf{r}\in\Pi$, and $A\subseteq\Pi$. When obvious from the context, we will omit the subscript $X$.
The semantics of $\mathcal{L}_{gather}$ is analogous to the one of $\mathcal{L}_{exp}$ in Definition 4.3, with the appropriate extensions to the interpretation function as already set.
Definition 4.35 (Approximate Point Gathering).
Any robot $\mathbf{r}$ strives to intersect their canonical position together with all other robots’ positions in a rendezvous area $u\in\text{P}_{rz(X)}$. Thus we say a system $\mathcal{I}$ is consistent with approximate point gathering, iff the following two conditions hold.
  • (Eventual Gathering): There is a rendezvous area where all robots eventually meet forever
    $\mathcal{I}\vDash\bigvee_{U\in\text{P}_{rz(X)}}\bigwedge_{\mathbf{r}\in\Pi}\Diamond\Box\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in U).$ (20)
  • (Starting Validity): If all robots are already contained in a rendezvous area initially, that is also the final rendezvous area
    $\begin{split}\mathcal{I}\vDash\bigwedge_{\begin{subarray}{c}U\in\text{P}_{rz(X)},\\ \vec{x}\in\text{P}_{g(X)}^{\Pi}\end{subarray}}\bigl{(}(\bigwedge_{r\in\Pi}init(&pos_{\mathbf{r}}(x_{\mathbf{r}})))\wedge(x_{\mathbf{r}}\in U)\bigr{)}\rightarrow\\ \bigl{(}\Diamond\Box\bigvee_{\vec{y}\in\text{P}_{g(X)}^{\Pi}}(\bigwedge_{\mathbf{r}\in\Pi}&pos_{\mathbf{r}}(y_{\mathbf{r}}))\wedge(y_{\mathbf{r}}\in U)\bigr{)}.\end{split}$ (21)
This strongly resembles the stabilizing agreement problem, epistemically characterized in [10]. It is defined by the following formulas, where $\mathcal{V}$ is the set of values to choose from, the atom $choose_{a}(v)$ denotes that agent $a$ chooses value $v$. $decide_{a}(v)$ is defined as $\Box choose_{a}(v)$, where $a,b$ are agents, i.e., robots with no dynamics as written in Example 2.3.
Definition 4.36 (Stabilizing Agreement [10]).
We say that a stable choice system is consistent with stabilizing agreement if (Agreement) and (Validity) hold.
  • (Agreement) : There is a value such that every agent decides on that value
    $\mathcal{I}\vDash\bigvee_{v\in\mathcal{V}}\bigwedge_{a\in\Pi}\Diamond decide_{a}(v)$ (22)
  • (Validity) : An agent can only choose a known initial value of some agent
    $\mathcal{I}\vDash\bigwedge_{v\in\mathcal{V}}\bigl{(}choose_{a}(v)\rightarrow K_{a}\bigvee_{b\in\Pi}init_{b}(v)\bigr{)}$ (23)
Even though the two conditions look syntactically equivalent, we have to do some work to show that stabilizing consensus implies approximate gathering. First observe that covering a rendezvous area can be seen as that robot choosing that area. In that sense, rendezvous areas are the objects the robots choose from. As each robot initially sits in exactly one area, that area becomes the problem input. The last obstacle is reducing the strong validity, as its called in the distributed computing literature (see for example [11]), to the weak validity of approximate gathering.
Theorem 4.37.
Consider a model $\mathcal{I}$ with an exploration space $X$, together with the language defined Definition 4.34, an arbitrary run $\sigma$ from $\mathcal{I}$. If $\mathcal{I}$ is consistent with stabilizing agreement, then it is also consistent with approximate point gathering.
Consider setting the set of possible choices $\mathcal{V}$ to be the set of possible rendezvous areas $\text{P}_{rz(X)}$, setting $choose_{\mathbf{r}}(U)\leftrightarrow\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in U)$ and $decide_{\mathbf{r}}(U)=\Box choose_{\mathbf{r}}(U)$, and setting $init_{\mathbf{r}}(U)\leftrightarrow\bigvee_{x\in\text{P}_{g(X)}}init(pos_{\mathbf{r}}(x))\wedge(x\in U)$.
Proof 4.38.
We start by substituting for $choose$ and $init$ as defined in the equations Equation 22 and Equation 23. For (Agreement) we immediately arrive at (Eventual Gathering):
$\mathcal{I}\vDash\bigvee_{U\in\text{P}_{rz(X)}}\bigwedge_{\mathbf{r}\in\Pi}\Diamond\Box\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in U).$ (24)
For (Starting Validity) we need to provide more reasoning, consider the following, where we already pre-pended the validity with a conjunction over all robots:
$\begin{split}\mathcal{I}\vDash\bigwedge_{\begin{subarray}{c}\mathbf{r}\in\Pi,\\ U\in\text{P}_{rz(X)}\end{subarray}}\bigl{(}(&\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in U))\rightarrow\\ (K_{\mathbf{r}}&\bigvee_{\mathbf{r}^{\prime}\in\Pi}\bigvee_{x\in\text{P}_{g(X)}}init(pos_{\mathbf{r}^{\prime}}(x))\wedge(x\in U))\bigr{)}.\end{split}$ (25)
As we need to prove an implication, we assume its antecedent and derive the consequent. Assuming
$\bigwedge_{\mathbf{r}\in\Pi}init(pos_{\mathbf{r}}(x_{\mathbf{r}}))\wedge(x_{\mathbf{r}}\in U),$ (26)
we want to derive
$\bigl{(}\Diamond\Box\bigvee_{\vec{y}\in\text{P}_{g(X)}^{\Pi}}(\bigwedge_{\mathbf{r}\in\Pi}pos_{\mathbf{r}}(y_{\mathbf{r}}))\wedge(y_{\mathbf{r}}\in U)\bigr{)},$ (27)
for any $U\in\text{P}_{rz(X)}$ and any $\vec{x}\in\text{P}_{g(X)}^{\Pi}$. Choose any $U$, $\vec{x}$ which satisfies Equation 26, meaning a starting position vector that intersects over $U$. Note that there is only one such $U$ as initial rendezvous areas are unique.
Observe that, as robots know their own initial positions, Equation 26 implies the antecedent in Equation 25 for all robots $\mathbf{r}$:
$\begin{split}&init(pos_{\mathbf{r}}(x_{\mathbf{r}}))\wedge(x_{\mathbf{r}}\in U)\implies\\ K_{\mathbf{r}}&init(pos_{\mathbf{r}}(x_{\mathbf{r}}))\wedge(x_{\mathbf{r}}\in U)\implies\\ K_{\mathbf{r}}\bigvee_{\mathbf{r}^{\prime}\in\Pi}&init(pos_{\mathbf{r}^{\prime}}(x_{\mathbf{r}^{\prime}}))\wedge(x_{\mathbf{r}^{\prime}}\in U)\implies\\ (K_{\mathbf{r}}\bigvee_{\mathbf{r}^{\prime}\in\Pi}\bigvee_{y\in\text{P}_{g(X)}}&init(pos_{\mathbf{r}^{\prime}}(y_{\mathbf{r}^{\prime}}))\wedge(y_{\mathbf{r}^{\prime}}\in U)).\end{split}$ (28)
Therefore the consequent in Equation 25 is true for only one choice of $U$ (the same that satisfies Equation 26) and false for all other $V\in\text{P}_{rz(X)}$.
By Equation 24 there exists some $V$, $t$ such that
$\mathcal{I},(\sigma,t)\vDash\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in V)$ (29)
holds for all $t^{\prime}>t$ and for all robots $r$. We see that Equation 29 is precisely the antecedent of Equation 25, and, as Equation 25 is a validity, it is satisfied by the fixed choice for $V$, $t$. Since the consequent in Equation 25 is true only for $U$ as derived in Equation 28 (but Equation 25 is a validity), Equation 29 has to be false for any $W\neq U$. But by the previous reasoning we know that Equation 29 is true for at least one $V$ for all $t^{\prime}>t$, implying that $U=V$ for all $t^{\prime}>t$. We can now derive Equation 27 from Equation 29:
$\begin{split}\bigvee_{x\in\text{P}_{g(X)}}&pos_{\mathbf{r}}(x)\wedge(x\in U)\\ \bigvee_{\vec{y}\in\text{P}_{g(X)}^{\Pi}}(\bigwedge_{\mathbf{r}\in\Pi}&pos_{\mathbf{r}}(\vec{y}_{r}))\wedge(x\in\vec{y}_{r})\bigr{)}\\ \bigl{(}\Diamond\Box\bigvee_{\vec{y}\in\text{P}_{g(X)}^{\Pi}}(\bigwedge_{\mathbf{r}\in\Pi}&pos_{\mathbf{r}}(\vec{y}_{r}))\wedge(x\in\vec{y}_{r})\bigr{)},\end{split}$ (30)
meaning the robots gather at the initially intersected rendezvous area forever.
This has non-trivial consequences: the second depth broadcaster condition identified in [10] immediately becomes a sufficient condition for approximate point gathering. As an application, the luminous robot model with full visibility is equivalent to the iterated immediate snapshot model [1] in the distributed computing view and therefore satisfies said second depth broadcaster condition. By the above reasoning, luminous with full visibility is immediately sufficient for approximate gathering.
5 Conclusions and future work
We developed a framework that integrates seemlessly control, distributed systems and (epistemic and temporal) logical aspects of multi-robot systems, with applications to exploration, surveillance and gathering tasks. This has multiple implications and holds the promise of multiple future developments. One of the most striking implications is that the classical luminous robot model [12] with full visibility is very strong in a distributed computing sense: not surprisingly, there is a whole research field on establishing and maintaining communication in multi robot systems. Some, such as [7, 8] have leveraged epistemic planning and action models in order to keep a live communication. As we have identified a sufficient condition for approximate gathering, we can directly judge the correctness of an action model, and therefore the correctness of a protocol, by its ability to reach a given logical specification.
As a follow-up of this work, we are considering extending the logic framework provided here by enriching the language with deontic modalities expressing robots’ preferences (or strategies) during protocol execution given their current local knowledge. We are planning to also extend our framework so as to integrate the geometrical interpretation [22] of epistemic frames and the recent advances [35] on sheaf theoretic and homological aspects of task solvability.
Appendix A Taxonomy of look-compute-move
There are several possible choices regarding the capacities of robots. The following Table 1 summarizes the most relevant variations found for robots under the look-compute-move abstraction of Definition 2.6. They are chosen in order to make a model mode or less powerful, with additional assumptions on their communication, observation and movement capacities.
Table 1: Partial glossary of LCM terminology.
Property Variant Description
Memory Oblivious No memory, all information is lost between rounds
Luminous Lights can be used to preserve information across rounds
Communication Silent No communication between robots
Luminous Information passed indirectly by retrieving lights during Look
Visibility Myopic Look retrieves a partial snapshot parametrized by view distance
Unlimited visibility Look retrieves a snapshot of the entire environment
Identity Anonymous No unique identifier
Indistinguishable No external identifier
Movement Rigid Move is always executed fully
Non-rigid Move can be interrupted, parametrized by a minimum distance
Dynamics Holonomic Move can be any arbitrary destination
Non-holonomic Move has to respect dynamic constraints, e.g. Dubins’ car
Orientation Oriented There is a global coordinate system that all can access
Disoriented Only different local coordinate systems
Chiral Different local coordinate systems, but with common north
Synchronicity ASYNC Scheduler, no shared notion of time
k-ASYNC Scheduler, the robots are at most $k$ rounds apart
SSYNC Scheduler, a subset of the robots activated at each round
FSYNC Scheduler, all robots are activated at each round